OpenStudy (anonymous):

The critical angle of refraction for calcite is 68.4 degrees when it forms a boundry with water. Using this information determine the speed of light in calcite.

OpenStudy (anonymous):

The critical angle of refraction for calcite is 68.4° when it forms a boundary with water. Use this information to determine the speed of light in calcite When a light ray travels from a more dense medium to a less dense medium, the angle between the light ray and the line perpendicular to the 2 surfaces increases. As the angle of incidence increases, the angle of refraction increases. At the critical angle of incidence, the angle of refraction = 90˚ This means the light ray is traveling along the line formed between the 2 surfaces. n1 * sin θi = n2 * sin θ r When θi = critical angle, θr = 90˚ Sin 90˚ = 1 n1 * sin θi = n2 * 1 sin θi = n2 / n1 sin 68.4˚ = n2 / n1 When determining the speed of light in a substance, n2 = 1, the index of refraction for a vacuum. sin 68.4 = 1 / n1 n1 = 1/ sin 68.4 The speed of light in calcite = n1 * c = c * 1/ sin 68.4 = 3.23 * 10^8 m/s

OpenStudy (anonymous):

i looked it up for you

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

NP #🌞🔑🌎